You carry out a reaction where 37.2 mL of a 0.779
Answer #1

moles of Ni(NO3)2 in 0.779M and 37.2 ml= 0.779*(mol/L)*37.2/1000 L=0.028979 moles

delH= -24.6 Kj/mol

1mole produces 24.6 Kj of heat

0.028979 moles produce 24.6*1000 joules*0.028979=712.8785 joules

enthalpy change -ve indicates heat is generated during the couse of reaction and q is+ve and q= 712.8785 joules